# From Continual Learning to SGD and Back: Better Rates for Continual Linear Models

### v2

@@ Appendix F Supplementary Material for the Extension Section (Section 6)
  • Proof.
    First, by Theorem 1.5.5 in Facchinei and Pang [17], $f_{m}$ is continuously differentiable and for every $\mathbf{w}\in\mathbb{R}^{d},m\in\left[{T}\right]$, $\nabla f_{m}(\mathbf{w})=\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})$. Plugging in $\nabla f_{\tau(t)}(\mathbf{w}_{t-1})$ into an appropriate SGD step, we get
@@
    and the second part of the lemma follows. In addition, $\forall\mathbf{x},\mathbf{w}\in\mathbb{R}^{d}$, we prove convexity by using a projection inequality (also from Theorem 1.5.5 in 17). That is,
@@
    where we used the non-expansiveness of $\mathbf{I}-\bm{\Pi}_{m}$ [Propositions 4.2, 4.8 in 3]. $\blacksquare$
@@
  • Proof.
    We have $\nabla f(\mathbf{w})=\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w})$. Hence, by Theorem 1.5.5 in Facchinei and Pang [17],
@@
    $=2f(\mathbf{w})+\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\bm{\Pi}_{\mathcal{K}}(\mathbf{w})-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
    $\geq 2f(\mathbf{w})-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\,.$
@@
    • Proof.
      The proof largely follows the same steps of Theorems 4.2 and 4.3. Let $\tau$ be any random ordering, $\mathbf{w}_{0}\in\mathbb{R}^{d}$ an initialization, and $\mathbf{w}_{1},\ldots,\mathbf{w}_{k}$ be the corresponding iterates produced by Scheme 4. By Reduction 3, these are exactly the (stochastic) gradient descent iterates produced when initializing at $\mathbf{w}_{0}$ and using a step size of $\eta=1$, on the $1$-smooth loss sequence $f_{\tau(1)},\ldots,f_{\tau(k)}$ defined by:
@@
[…]
    which proves the without-replacement case and thus completes the proof. $\blacksquare$
@@
[…]
@@
    which completes the proof. $\blacksquare$
@@
  • Proof.
    Our proof here mostly follows the proof of Lemma E.6. Recall that from Eq. (8), any $\beta$-smooth realizable function $h:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ holds that $\left|h(\tilde{\mathbf{w}})-h(\mathbf{w})\right|\leq h(\mathbf{w})+\beta\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$. Denote $f_{t}\triangleq f(\cdot;i_{t})$ for all $t\in\left\{0,.\hskip 0.8pt\!.\hskip 0.8pt\!.,T\right\}$. Now, by the standard stability $\iff$ generalization argument [61, 25], and denoting by $\mathbf{w}_{\tau}^{(i)}$ the SGD iterate after $\tau$ steps on the training set where the $i$th example was resampled as $j_{i}$:
@@
    Next, we bound $\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$. By the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ [see Lemma 3.6 in 25]:
@@
    and the result follows. $\blacksquare$

### v3

@@ Appendix F Supplementary Material for the Extension Section (Section 6)
Proof F.1.
First, by Theorem 1.5.5 in Facchinei and Pang (2003), $f_{m}$ is continuously differentiable and for every $\mathbf{w}\in\mathbb{R}^{d},m\in\left[{T}\right]$, $\nabla f_{m}(\mathbf{w})=\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})$. Plugging in $\nabla f_{\tau(t)}(\mathbf{w}_{t-1})$ into an appropriate SGD step, we get
@@
and the second part of the lemma follows. In addition, $\forall\mathbf{x},\mathbf{w}\in\mathbb{R}^{d}$, we prove convexity by using a projection inequality (also from Theorem 1.5.5 in Facchinei and Pang, 2003). That is,
@@
where we used the non-expansiveness of $\mathbf{I}-\bm{\Pi}_{m}$ (Propositions 4.2, 4.8 in Bauschke et al., 2011).
@@
Proof F.3.
We already established that $\nabla f(\mathbf{w})=\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w})$. Combining this with simple algebra, we obtain,
@@
$=2f(\mathbf{w})+\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\bm{\Pi}_{\mathcal{K}}(\mathbf{w})-\mathbf{z}\right\rangle$
$=2f(\mathbf{w})+\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\bm{\Pi}_{\mathcal{K}}(\mathbf{w})-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\,.$
By Theorem 1.5.5 (b) in Facchinei and Pang (2003), we have that
$\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\bm{\Pi}_{\mathcal{K}}(\mathbf{w})-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\geq 0\,$,
and finally we get,
$\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $\geq 2f(\mathbf{w})-\left\langle\mathbf{w}-\bm{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-\bm{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\,.$
@@
Proof F.4.
The proof largely follows the same steps of Theorems 4.3 and 4.4. Let $\tau$ be any random ordering, $\mathbf{w}_{0}\in\mathbb{R}^{d}$ an initialization, and $\mathbf{w}_{1},\ldots,\mathbf{w}_{k}$ be the corresponding iterates produced by Scheme 4. By Reduction 3, these are exactly the (stochastic) gradient descent iterates produced when initializing at $\mathbf{w}_{0}$ and using a step size of $\eta=1$, on the $1$-smooth loss sequence $f_{\tau(1)},\ldots,f_{\tau(k)}$ defined by:
@@
$\leq\left({\frac{7k}{T}+\frac{\frac{e}{2}(T-k)}{T}}\right)\frac{1}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,$,
which proves the without-replacement case and thus completes the proof.
@@
$\leq\left({\frac{e}{\sqrt[4]{k}}+\frac{4}{k}}\right)\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}\leq\frac{7\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}.$
@@
which completes the proof.
@@
Proof F.7.
Our proof here mostly follows the proof of Lemma E.11. Recall that from Eq. (8), any $\beta$-smooth realizable function $h:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ holds that $\left|h(\tilde{\mathbf{w}})-h(\mathbf{w})\right|\leq h(\mathbf{w})+\beta\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$. Denote $f_{t}\triangleq f(\cdot;i_{t})$ for all $t\in\left\{0,.\thinspace\!.\thinspace\!.,T\right\}$. Now, by the standard stability $\iff$ generalization argument (Shalev-Shwartz et al., 2010; Hardt et al., 2016), and denoting by $\mathbf{w}_{\tau}^{(i)}$ the SGD iterate after $\tau$ steps on the training set where the $i$th example was resampled as $j_{i}$:
@@
Next, we bound $\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$. By the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in Hardt et al., 2016):
@@
and the result follows.
Appendix G Supplementary Material for the Discussion Section (Section 7)
Claim 1 (Average-Norm Universal Rate for With-Replacement Random Ordering).
$\mathbb{E}_{\tau}\!\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $\leq\frac{2\left\|{\mathbf{w}_{\star}}\right\|^{2}\bar{R}}{\sqrt[4]{k}}\,,\quad\quad\mathbb{E}_{\tau}\!\left[F_{\tau}(k)\right]\leq\frac{5\left\|{\mathbf{w}_{\star}}\right\|^{2}\bar{R}}{\sqrt[4]{k-1}}\,$,
where $\bar{R}=\sum_{m=1}^{M}\left\|{\mathbf{X}_{m}}\right\|^{2}/T$.
Proof sketch. Taking the non-worst case bound from Lemma 3.1, we have $\mathcal{L}_{m}(w)\leq\alpha_{m}f_{m}(w)$ for $\alpha_{m}=\|\mathbf{X}_{m}\|^{2}$. Then in the proof of Theorem 4.3, $\mathcal{L}(w)\leq\frac{A}{2T}\sum\frac{\alpha_{m}}{A}f_{m}(w)$, where $A=\sum\alpha_{m}$, and we may apply Theorem 5.1 (which supports arbitrary distributions $\mathcal{D}$, see Setup 1) with the distribution given by $\Pr_{\mathcal{D}}(i)=\alpha_{i}/A$. Finally, we have $\bar{R}=A/T$.